Does y = sin (kx − ωt) satisfies the wave equation? Justify your answer
1
Expert's answer
2019-05-13T10:29:33-0400

The wave equation:


"v^2 \frac{\partial^2 y}{\partial x^2}=\frac{\partial^2 y}{\partial t^2}"

We have:


"\frac{\partial y}{\partial x}=k \cos{(kx − ωt)}"

"\frac{\partial^2 y}{\partial x^2}=-k^2 \sin (kx − ωt)"

"\frac{\partial y}{\partial t}=-\omega \cos{(kx − ωt)}"

"\frac{\partial^2 y}{\partial t^2}=-\omega^2 \sin (kx − ωt)"

Thus,


"v^2 (-k^2 \sin (kx − ωt) )=-\omega^2 \sin (kx − ωt)"

We have the wave equation with


"v=\frac{\omega}{k}"